# LiCoMemory: Lightweight and Cognitive Agentic Memory for Efficient Long-Term Reasoning

### v1

@@ Abstract
Zhengjun Huang1, Zhoujin Tian1, Qintian Guo1, Fangyuan Zhang2, Yingli Zhou3, Di Jiang4, Xiaofang Zhou1 1The Hong Kong University of Science and Technology; 2Huawei Hong Kong Research Center, Hong Kong; 3The Chinese University of Hong Kong, Shenzhen; 4WeBank, Shenzhen zhuangff@cse.ust.hk, ztianaf@cse.ust.hk, qtguo@ust.hk, zhang.fangyuan@huawei.com yinglizhou@link.cuhk.edu.cn, dijiang@webank.com, zxf@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable advancements across a wide range of language understanding and generation tasks Achiam et al. (2023); Bai et al. (2023) and are increasingly evolving into personalized assistants with enhanced contextual reasoning capabilities Li et al. (2024). Despite their strong generalizing and reasoning abilities, LLMs remain constrained by a critical short-term memory limitation: the finite context window. Information beyond the context window cannot be effectively preserved or recalled, leading to degraded reasoning capability and reduced response accuracy in long-term conversational scenarios Hatalis et al. (2023).
@@
To resolve these issues, early attempts to enhance the long-term memory of agents commonly adopt Retrieval-Augmented Generation (RAG) architectures, which leverage conversation history as an external knowledge source and retrieve contextually relevant information to support response generation Singh et al. (2025); Zhang et al. (2025b). Graph-based RAG further extends this paradigm by structuring conversational content into relational graphs using heuristic rules, thereby capturing semantic dependencies among historical records and improving cross-session reasoning Zhu et al. (2025). Despite their effectiveness in retrieval, these approaches often treat memory as a static component, neglecting the inherently dynamic nature of human–agent interactions, and thus lacking mechanisms for reorganization of accumulated knowledge. Moreover, such predefined retrieval and linking strategies can lead to information loss and hinder adaptive memory evolution Li et al. (2025a). Recent studies have thus shifted toward dynamic agent memory frameworks, which extend beyond traditional RAG by modeling the evolving nature of conversational data and adaptive retrieval. Mem0 Chhikara et al. (2025) initiates this shift by introducing explicit memory operations that allow agents to manage and revise stored knowledge as conversations evolve. Building on this, MIRIX Wang and Chen (2025) enhances retrieval organization through multi-granularity memory indexing and relevance fusion, improving contextual alignment and reducing redundancy in long-term reasoning. Further, Zep Rasmussen et al. represents memory as a graph to capture relational dependencies among dialogue events, promoting interpretability but at the cost of high graph-construction overhead and retrieval latency.
@@
While these frameworks have made progress toward structured conversational memory, some key challenges still remain. (1) Coupled and redundant graph structures: Existing graph-based memory systems often intertwine semantic content with relational topology, leading to heavy, redundant, and inflexible graph representations that are difficult to adapt to dynamic human–agent interactions Chen et al. (2025). (2) Scattered and unstructured retrieval: Owing to flat architectures and unguided retrieval mechanisms, conventional memory pipelines frequently return fragmented or contextually inconsistent information, which undermines reasoning coherence and produces semantically diluted responses Xiang et al. (2025). (3) Slow update and inference: Large, monolithic graph structures incur substantial computational overhead during incremental updates and inference, limiting their applicability in real-time interaction settings. For instance, GraphRAG Edge et al. (2024a) requires up to 20 minutes for graph and community construction per dialogue and over 2 minutes of query latency Zhang et al. .
@@
To address these challenges, we propose LiCoMemory, an end-to-end agentic memory framework that enables real-time updating and retrieval. At its core, LiCoMemory introduces CogniGraph, a lightweight and semantically aware hierarchical graph that redefines the knowledge graph as a semantic indexing layer rather than a static repository. By using graph topology as a structural scaffold instead of embedding extensive content within nodes and edges, CogniGraph indexes and organizes knowledge while linking relational structures to their original textual sources for precise and context-aware reasoning. During inference, LiCoMemory integrates a unified reranking mechanism that jointly considers semantic similarity, hierarchical structure, and temporal relevance to achieve accurate and structured retrieval. Experimental results demonstrate that LiCoMemory achieves up to 23% improvement in accuracy over the second-best baseline on established long-term dialogue benchmarks (LoCoMo Maharana et al. (2024) and LongMemEval Wu et al. (2024)), particularly on multi-session and temporal reasoning subsets, while significantly reducing input tokens and response latency, underscoring its efficiency.
@@ 2.1 Retrieval-Augmented Generation (RAG)
RAG has emerged as a foundational framework for augmenting LLMs with external memory Fan et al. (2024). A typical RAG pipeline first splits prior interactions or knowledge as segments, then retrieves relevant segments during inference, and provides both the user query and retrieved content to the language model for grounded generation Zhou et al. (2025). Due to the highly entangled and semantically redundant nature of conversational data, conventional RAG often retrieves overlapping or loosely related content, resulting in fragmented context and limited reasoning continuity Li et al. (2025b). To address these limitations, recent research has introduced Graph-based RAG to agent frameworks, where the historical interactions are pre-organized as a graph structure to improve retrieval and reasoning efficiency Edge et al. (2024b). Beyond flat graph representations, several frameworks further organize memory knowledge hierarchically, such as tree-based structures Zhang et al. (2025a) and community-aware architectures Wang et al. (2025). These structured organizations capture cross-session dependencies, reduce redundancy, and enhance contextual coherence, paving the way for more structured and adaptive memory systems.
@@ 2.2 Agent Memory Augmentation
While LLM-based agents demonstrate remarkable generative capabilities, they remain constrained by limited context windows and the absence of persistent memory, often resulting in inconsistent behavior across extended interactions. Agent memory augmentation therefore emerges as a promising direction to address this limitation, aiming to equip conversational agents with external memory systems that support long-term information retention, retrieval, and reasoning. LoCoMo Maharana et al. (2024) introduce a RAG-style conversational framework capable of maintaining multi-session dialogue continuity through chunk-based retrieval and coherence-aware evaluation. A-MEM Xu et al. (2025) advances the concept of agentic memory beyond passive long-term storage by introducing a self-organizing, dynamically evolving memory architecture that autonomously constructs, links, and refines knowledge representations. LongMemEval Wu et al. (2024) further advances this line of research by proposing a dedicated long-term memory model and benchmark for evaluating temporal reasoning, knowledge updating, and cross-session consistency. Mem0 Chhikara et al. (2025) employs a scalable two-phase architecture (extraction and update) that dynamically stores and retrieves salient facts using a vector database, while also offering a graph-based variant that better supports long term memory maintenance and retrieval. More recently, Zep Rasmussen et al. structures agent memory into knowledge graphs, improving retrieval relevance but suffering from high construction overhead and retrieval latency.
@@ 3.2 Query Processing and Integrated Rerank
$w(\Delta\tau)=\exp\!\Big[-\Big(\frac{\Delta\tau}{\hat{\tau}}\Big)^{k}\Big],\quad 0<k<1,$
where $\Delta\tau$ denotes the time gap between the current query and the triple’s timestamp, and $\hat{\tau}$ is the median of time gaps across retrieved triples, allowing the decay to adapt to the temporal scale of each query. Finally, the overall relevance score of each triple is defined as $R(t)=S_{\text{sem}}\times w(\Delta\tau)$, which jointly captures semantic coherence and temporal recency. Because $w(\Delta\tau)\in(0,1]$, the time dimension modulates rather than dominates the semantic signal, ensuring that highly relevant but moderately older information remains retrievable. This formulation achieves a balanced trade-off between temporal adaptivity and semantic precision. The final output is a structured prompt that consolidates highly relevant session summaries, top-ranked triples, and their corresponding original dialogue chunks from which the triples were extracted. This structured representation provides the language model with a coherent and fine-grained view of the relevant knowledge, facilitating precise reasoning across different levels of granularity and enabling robust cross-session understanding.
@@ 4 Experiments
In this section we evaluate LiCoMemory on real-world datasets to assess its performance. In particular, we aim to answer the following research questions: Q1: How effective is LiCoMemory compared with existing memory paradigms? Q2: How does the components of our system affect the final result? Q3: How does LiCoMemory perform in real-time practical scenarios?
Table 1: Evaluation on long-term memory QA benchmarks utilizing different language models. The best and second-best results are highlighted. $T_{R}$ stands for query latency.
@@ 4.1 Experimental Setup
Dataset. The performance of LiCoMemory was evaluated on two public long-term memory benchmarks: LongMemEval and LoCoMo. LongMemEval Wu et al. (2024) is a comprehensive benchmark for evaluating long-term memory in conversational agents, consists of 500 questions across six types: single-session user (S.S.U.), single-session assistant (S.S.A.), single-session preference (S.S.P.), multi-session, temporal reasoning, and knowledge update. LoCoMo Maharana et al. (2024) focuses on extremely long multi-session dialogues, containing 1,986 questions in five distinct categories: single-hop, multi-hop, temporal, open-domain and adversial reasoning.
Metrics. We evaluate different methods from two aspects: response quality and efficiency. Following previous work Wu et al. (2024), we assess response quality using Accuracy (Acc.), Recall (Rec.), and F1 score. For accuracy, we adopt the same evaluation protocol in LongMemEval Wu et al. (2024), where a large language model makes binary assessments of answer correctness. Recall is defined as the proportion of ground-truth targets correctly retrieved within the top-15 items Xu et al. (2025). And F1 score is reported as the harmonic mean of accuracy and recall. Efficiency is measured by token consumption and query latency Chhikara et al. (2025), which respectively represent the total number of tokens consumed by the LLM during query processing and the time required to complete the entire response generation.
Baselines. We compare LiCoMemory with four well-established baselines. LoCoMo Maharana et al. (2024) segments entire dialogues into textual chunks and retrieves the most relevant chunks to answer questions; Zep Rasmussen et al. employs structured memory access strategies to address complex, temporally extended queries; Mem0 Chhikara et al. (2025) performs explicit in-context memory operations for interactive agent environments; A-MEM Xu et al. (2025) introduces a dynamically evolving memory architecture to link and refine knowledge representations.
Implementation Details. All methods are evaluated under the same settings. Llama-3.1-70B-Instruct-Turbo serves as the primary LLM for memory management, while BGE-M3 is adopted for text embedding to support retrieval. (citation) During answer generation, Llama-3.1-70B-Instruct-Turbo and GPT-4o-mini are utilized as generation models. The number of retrieved memory units is set to 15. All experiments are conducted on NVIDIA 80G A100 GPUs.
@@ 4.2 Main Results(RQ1)
Table 1 presents a comprehensive comparison of LiCoMemory with representative memory frameworks across two long-term dialogue benchmarks using different backbone language models. Across both Llama-3.1-70B-Instruct-Turbo and GPT-4o-mini, LiCoMemory consistently achieves the highest accuracy and recall while maintaining the lowest or near-lowest query latency, demonstrating great robustness. Specifically, on LongMemEval, it surpasses the second-best baseline (Zep) by 9.0% in accuracy and 9.7% in recall with Llama-3.1-70B, and by 11.2% and 5.3% respectively with GPT-4o-mini. Similarly, on LoCoMo, LiCoMemory outperforms Mem0 by 9.8% in accuracy and 7.5% in recall on Llama-3.1-70B, and achieves a 12.5% and 5.8% gain under GPT-4o-mini. The observed improvement in QA performance underscores the effectiveness of the proposed CogniGraph structure, while the reduced latency and retrieval volume further highlight the benefits of its precise retrieval mechanism and lightweight, efficiency-oriented graph design.
To further analyze how LiCoMemory outperforms other frameworks on long-term dialogue benchmarks, we provide a detailed breakdown of its performance across subsets with different focuses using GPT-4o-mini as the generation model, as shown in Figure 4. As illustrated in the left chart, LiCoMemory consistently surpasses the second-best baseline (Mem0) across all subsets of LoCoMo, with a large gain observed in the Temporal-Reasoning subset, where accuracy improves by 20.7%. A smaller improvement is observed on the Adversarial subset, likely due to occasional false positives arising when the correct entries are not retrieved, as reflected by the recall results in Table 1. On the LongMemEval benchmark, LiCoMemory achieves substantial gains on the Multi-Session (26.6%) and Temporal Reasoning (15.9%) subsets, highlighting the effectiveness of our CogniGraph structure and unified reranking mechanism in capturing temporal and cross-session dependencies.
@@
4.3 Ablation Study(RQ2)
Figure 5: Performance breakdown of LiCoMemory and baselines on LoCoMo and LongmemEval.
In this section, we investigate how each major component of LiCoMemory contributes to different aspects of its capability. We individually disable key modules and evaluate their effects on both Accuracy and Recall across several representative scenarios within the LongMemEval benchmark. As shown in Figure 5, removing different components results in varying degrees of performance decrease.
w/o Structured retrieval. We remove the structured retrieval mechanism of CogniGraph, performing retrieval solely over extracted triples without considering the hierarchical organization of sessions or entity–relation links. As shown in Figure 5, this simplification causes the most severe degradation among all variants, dropping from 73.8/80.6 to 51.6/56.4 on the full dataset, and especially in the multi-session setting, where performance plunges from 64.7/73.5 to 40.6/48.2. Without structured aggregation under shared contextual summaries or temporal anchors, retrieval becomes fragmented and factual grounding weakens. These results highlight that structured retrieval is essential for maintaining cross-session coherence, linking fine-grained facts to their contextual origins, and enabling robust long-term reasoning.
w/o Temporal awareness. We remove the temporal weighting mechanism in re-ranking process and perform retrieval solely based on semantic similarity, without considering timestamp-based decay and temporal information reasoning. As shown in Figure 5, in the temporal reasoning setting, accuracy and recall drop sharply from 71.4/81.3 to 48.9/53.5, and in knowledge update, performance decreases from 80.8/85.1 to 51.9/57.0. In these time-sensitive scenarios, this ablation leads to the largest performance degradation among all variants, as the model overemphasizes semantically similar but outdated triples, causing obsolete or conflicting information to be retrieved. These findings underscore the importance of temporal awareness in preserving chronological coherence and enabling reliable reasoning over dynamically evolving knowledge in long-term dialogues.
w/o Summary. We remove the summary-level weighting from ranking process, forcing the model to rely solely on fine-grained triples for relevance estimation. As shown in Figure 5, this modification consistently degrades performance across all settings, dropping from 73.8/80.6 to 61.4/69.5 on the full dataset and from 64.7/73.5 to 42.9/51.3 in the multi-session case. Without the guidance of high-level summaries, retrieval becomes overly local and miss broader contextual cues that link related sessions and entities. These observations indicate that summary plays a crucial role in abstracting long-term context and steering the retrieval process toward semantically coherent evidence.
4.4 Detailed Case Study(RQ3)
Following the case discussed in Section 3.3, we now conduct a detailed evaluation of how LiCoMemory performs in a practical case, where a user engages with the agent and demands real-time update and retrieval from LiCoMemory. For this experiment, we use Llama-3.1-70B-Instruct-Turbo as the backbone model and perform a chunk by chunk insertion on the LongmemEval dataset to simulate the real life cases. During the insertion of the context dialogues, we do not request retrieval from CogniGraph but instead only insert the dialogue chunks, and perform retrieval for the query. We collect the data and present it as follows. $K_{G}$ denotes the average token consumption per session during the context processing stage, reflecting the system’s efficiency in summarizing and structuring dialogue history. $K_{R}$ represents the number of input tokens in the retrieval prompt, indicating the compactness of the information fed into the language model during response generation. $T_{G}$ measures the latency per session for context processing, while $T_{R}$ quantifies the query latency at inference time.As shown, LiCoMemory achieves an accuracy of 67.4%, with only 1.8% of accuracy drop compared to the static scenario, demonstrating strong robustness. Meanwhile LiCoMemory computes with modest token usage (13.52k for $K_{G}$ and 1.22k for $K_{R}$) and low latency (5.21 s for $T_{G}$ and 2.63 s for $T_{R}$), demonstrating its capability to maintain both retrieval accuracy and real-time responsiveness in dynamic conversational scenarios.

### v2

@@ Abstract
Zhengjun Huang1, Zhoujin Tian1, Qintian Guo1, Fangyuan Zhang2, Yingli Zhou3, Di Jiang4, Zeying Xie4, Xiaofang Zhou1 1The Hong Kong University of Science and Technology; 2Huawei Hong Kong Research Center, Hong Kong; 3The Chinese University of Hong Kong, Shenzhen; 4WeBank Co., Ltd, Shenzhen zhuangff@cse.ust.hk, ztianaf@cse.ust.hk, qtguo@ust.hk, zhang.fangyuan@huawei.com yinglizhou@link.cuhk.edu.cn, dijiang@connect.ust.hk, zenxie@webank.com, zxf@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable advancements across a wide range of language understanding and generation tasks gpt; qwen and are increasingly evolving into personalized assistants with enhanced contextual reasoning capabilities llmagents. Despite their strong generalizing and reasoning abilities, LLMs remain constrained by a critical short-term memory limitation: the finite context window. Information beyond the context window cannot be effectively preserved or recalled, leading to degraded reasoning capability and reduced response accuracy in long-term conversational scenarios needmemory.
@@
To resolve these issues, early attempts to enhance the long-term memory of agents commonly adopt Retrieval-Augmented Generation (RAG) architectures, which leverage conversation history as an external knowledge source and retrieve contextually relevant information to support response generation ragagent; ragsurvey. Graph-based RAG further extends this paradigm by structuring conversational content into relational graphs using heuristic rules, thereby capturing semantic dependencies among historical records and improving cross-session reasoning graph. Despite their effectiveness in retrieval, these approaches often treat memory as a static component, neglecting the inherently dynamic nature of human–agent interactions, and thus lacking mechanisms for reorganization of accumulated knowledge. Moreover, such predefined retrieval and linking strategies can lead to information loss and hinder adaptive memory evolution ragagentflaw. Recent studies have thus shifted toward dynamic agent memory frameworks, which extend beyond traditional RAG by modeling the evolving nature of conversational data and adaptive retrieval. Mem0 mem0 initiates this shift by introducing explicit memory operations that allow agents to manage and revise stored knowledge as conversations evolve. Building on this, MIRIX mirix enhances retrieval organization through multi-granularity memory indexing and relevance fusion, improving contextual alignment and reducing redundancy in long-term reasoning. Further, Zep zep represents memory as a graph to capture relational dependencies among dialogue events, promoting interpretability but at the cost of high graph-construction overhead and retrieval latency.
@@
While these frameworks have made progress toward structured conversational memory, some key challenges still remain. (1) Coupled and redundant graph structures: Existing graph-based memory systems often intertwine semantic content with relational topology, leading to heavy, redundant, and inflexible graph representations that are difficult to adapt to dynamic human–agent interactions prob1. (2) Scattered and unstructured retrieval: Owing to flat architectures and unguided retrieval mechanisms, conventional memory pipelines frequently return fragmented or contextually inconsistent information, which undermines reasoning coherence and produces semantically diluted responses prob2. (3) Slow update and inference: Large, monolithic graph structures incur substantial computational overhead during incremental updates and inference, limiting their applicability in real-time interaction settings. For instance, GraphRAG graphrag requires up to 20 minutes for graph and community construction per dialogue and over 2 minutes of query latency workshop1.
@@
To address these challenges, we propose LiCoMemory, an end-to-end agentic memory framework that enables real-time updating and retrieval. At its core, LiCoMemory introduces CogniGraph, a lightweight and semantically aware hierarchical graph that redefines the knowledge graph as a semantic indexing layer rather than a static repository. By using graph topology as a structural scaffold instead of embedding extensive content within nodes and edges, CogniGraph indexes and organizes knowledge while linking relational structures to their original textual sources for precise and context-aware reasoning. During inference, LiCoMemory integrates a unified reranking mechanism that jointly considers semantic similarity, hierarchical structure, and temporal relevance to achieve accurate and structured retrieval. Experimental results demonstrate that LiCoMemory achieves up to 23% improvement in accuracy over the second-best baseline on established long-term dialogue benchmarks (LoCoMo locomo and LongMemEval longmem), particularly on multi-session and temporal reasoning subsets, while significantly reducing input tokens and response latency, underscoring its efficiency.
@@ 2.1 Retrieval-Augmented Generation (RAG)
RAG has emerged as a foundational framework for augmenting LLMs with external memory rag1. A typical RAG pipeline first splits prior interactions or knowledge as segments, then retrieves relevant segments during inference, and provides both the user query and retrieved content to the language model for grounded generation depth. Due to the highly entangled and semantically redundant nature of conversational data, conventional RAG often retrieves overlapping or loosely related content, resulting in fragmented context and limited reasoning continuity ragprob. To address these limitations, recent research has introduced Graph-based RAG to agent frameworks, where the historical interactions are pre-organized as a graph structure to improve retrieval and reasoning efficiency gr1. Beyond flat graph representations, several frameworks further organize memory knowledge hierarchically, such as tree-based structures erarag and community-aware architectures archrag. These structured organizations capture cross-session dependencies, reduce redundancy, and enhance contextual coherence, paving the way for more structured and adaptive memory systems.
@@ 2.2 Agent Memory Augmentation
While LLM-based agents demonstrate remarkable generative capabilities, they remain constrained by limited context windows and the absence of persistent memory, often resulting in inconsistent behavior across extended interactions. Agent memory augmentation therefore emerges as a promising direction to address this limitation, aiming to equip conversational agents with external memory systems that support long-term information retention, retrieval, and reasoning. LoCoMo locomo introduce a RAG-style conversational framework capable of maintaining multi-session dialogue continuity through chunk-based retrieval and coherence-aware evaluation. A-MEM amem advances the concept of agentic memory beyond passive long-term storage by introducing a self-organizing, dynamically evolving memory architecture that autonomously constructs, links, and refines knowledge representations. LongMemEval longmem further advances this line of research by proposing a dedicated long-term memory model and benchmark for evaluating temporal reasoning, knowledge updating, and cross-session consistency. Mem0 mem0 employs a scalable two-phase architecture (extraction and update) that dynamically stores and retrieves salient facts using a vector database, while also offering a graph-based variant that better supports long term memory maintenance and retrieval. More recently, Zep zep structures agent memory into knowledge graphs, improving retrieval relevance but suffering from high construction overhead and retrieval latency.
@@ 3.2 Query Processing and Integrated Rerank
$w(\Delta\tau)=\exp\!\Big[-\Big(\frac{\Delta\tau}{\hat{\tau}}\Big)^{t_{k}}\Big],\quad 0<t_{k}<1,$
where $\Delta\tau$ denotes the time gap between the current query and the triple’s timestamp, and $\hat{\tau}$ is the median of time gaps across retrieved triples, allowing the decay to adapt to the temporal scale of each query. $t_{k}$ denotes the decay coefficient, where a larger value imposes a stronger penalty on temporally distant information. Finally, the overall relevance score of each triple is defined as $R(t)=S_{\text{sem}}\times w(\Delta\tau)$, which jointly captures semantic coherence and temporal recency. Because $w(\Delta\tau)\in(0,1]$, the time dimension modulates rather than dominates the semantic signal, ensuring that highly relevant but moderately older information remains retrievable. This formulation achieves a balanced trade-off between temporal adaptivity and semantic precision. The final output is a structured prompt that consolidates highly relevant session summaries, top-ranked triples, and their corresponding original dialogue chunks from which the triples were extracted. This structured representation provides the language model with a coherent and fine-grained view of the relevant knowledge, facilitating precise reasoning across different levels of granularity and enabling robust cross-session understanding.
@@ 4 Experiments
In this section we evaluate LiCoMemory on real-world datasets to assess its performance. In particular, we aim to answer the following research questions: Q1: How effective is LiCoMemory compared with existing memory paradigms? Q2: How does LiCoMemory perform in real-time practical scenarios? Q3: How does the components of our system affect the final result?
Table 1: Evaluation on long-term memory QA benchmarks utilizing different language models. The best and second-best results are highlighted. $T_{R}$ stands for query latency while $K_{R}$ stands for prompt token consumption. All results are averaged over 5 runs and we report mean $\pm$ standard deviation.
@@ 4.1 Experimental Setup
Dataset. The performance of LiCoMemory was evaluated on two public long-term memory benchmarks: LongMemEval and LoCoMo. LongMemEval longmem is a comprehensive benchmark for evaluating long-term memory in conversational agents, consists of 500 questions across six types: single-session user (S.S.U.), single-session assistant (S.S.A.), single-session preference (S.S.P.), multi-session, temporal reasoning, and knowledge update. LoCoMo locomo focuses on extremely long multi-session dialogues, containing 1,986 questions in five distinct categories: single-hop, multi-hop, temporal, open-domain and adversial reasoning. Detailed statistics can be found in Appendix A.1.
Metrics. We evaluate different methods from two perspectives: response quality and efficiency. Following prior work longmem, response quality is measured using Accuracy (Acc.) and Recall (Rec.), with the evaluation prompts detailed in Appendix A.3. Accuracy is assessed using the LLM-as-a-Judge protocol from LongMemEval longmem, in which a large language model performs binary judgments of answer correctness. As partial credit is not permitted, this metric provides a more faithful, human-aligned estimate of retrieval quality. For the Adversarial subset of the LoCoMo dataset, where ground-truth answers are unavailable, all responses are labeled as “Context insufficient to answer.” Recall is defined as the proportion of ground-truth targets retrieved within the top-15 items. Efficiency is evaluated in terms of token consumption and query latency mem0, corresponding to the total number of tokens used by the LLM during query processing and the time required to generate a complete response.
Baselines. We compare LiCoMemory with several well-established baselines, including LoCoMo locomo, Zep zep, Mem0(Mem0gmem0, A-Mem amem, Memorybank Memorybank, and MemOS memos. Detailed introduction is listed in Appendix A.2.
Implementation Details. All methods are evaluated under the same settings. Llama-3-8B-Instruct serves as the primary LLM for memory construction, while BGE-M3 is adopted for text embedding to support retrieval. During answer generation, Llama-3.1-70B-Instruct-Turbo and GPT-4o-mini are utilized as generation models. All experiments are conducted on NVIDIA 80G A100 GPUs. We report results averaged over 5 independent runs to ensure reliable performance and runtime measurements. The number of retrieved memory units ($k$) is set to 15 and the decay coefficient ($t_{k}$) is set to 0.1 with details shown in Appendix A.4.
@@ 4.2 Main Results(RQ1)
Table 1 presents a comprehensive comparison of LiCoMemory with representative memory frameworks across two long-term dialogue benchmarks using different backbone language models. Across both Llama-3.1-70B-Instruct-Turbo and GPT-4o-mini, LiCoMemory consistently achieves the highest accuracy and recall while maintaining the lowest or near-lowest query latency. Moreover, its performance remains stable across multiple runs, demonstrating strong robustness. Specifically, on LongMemEval, it surpasses the second-best baseline by 9.0% in accuracy and 9.3% in recall with Llama-3.1-70B, and by 9.0% and 5.3% respectively with GPT-4o-mini. Similarly, on LoCoMo, LiCoMemory outperforms Mem0g by 7.5% in accuracy and 5.2% in recall on Llama-3.1-70B, and achieves a 8.9% and 4.95% gain under GPT-4o-mini. The observed improvements in QA performance demonstrate the effectiveness of the proposed CogniGraph structure. In addition, the reductions in retrieval latency and retrieval volume further substantiate the advantages of its precise retrieval mechanism and lightweight, efficiency-oriented graph design. Notably, the most pronounced performance gains are observed on the LoCoMo benchmark, where retrieval latency is reduced by 10% and token consumption by 45% compared to the second-best baseline (Mem0).
To further analyze how LiCoMemory outperforms other frameworks on long-term dialogue benchmarks, we provide a detailed breakdown of its performance across subsets with different focuses using GPT-4o-mini as the generation model, as shown in Figure 4. As illustrated in the left chart, LiCoMemory consistently surpasses the second-best baseline (MemOS) across all subsets of LoCoMo, with a large gain observed in the Temporal-Reasoning subset, where accuracy improves by 19.2%. A smaller improvement is observed on the Adversarial subset, likely due to occasional false positives arising when the correct entries are not retrieved, as reflected by the recall results in Table 1. On the LongMemEval benchmark, LiCoMemory achieves substantial gains on the Multi-Session (26.6%) and Temporal Reasoning (15.9%) subsets compared to the second best baseline (Mem0), highlighting the effectiveness of our CogniGraph structure and unified reranking mechanism in capturing temporal and cross-session dependencies.
@@
Figure 5: Ablation study of LiCoMemory on LongmemEval and subset breakdown.
4.3 Real-Time Performance(RQ2)
Following the discussion in Section 3.3, we evaluate LiCoMemory in a practical interactive setting where the agent must support real-time updates and retrieval. Using Llama-3.1-70B-Instruct-Turbo as the backbone, we perform chunk-by-chunk insertion on the LoCoMo dataset to emulate real-world conversational flows. During context ingestion, we only insert dialogue chunks without triggering retrieval, and issue queries after all insertions are completed. As shown in Table 2, LiCoMemory maintains leading accuracy with minimal degradation from static to real-time insertion, while achieving the lowest token usage and latency in both context processing ($K_{G}$, $T_{G}$) and querying ($K_{R}$, $T_{R}$). With the support of CogniGraph, LiCoMemory further reduces construction token cost by over 3 times and construction latency by more than an order of magnitude, without compromising retrieval quality. These results highlight both the efficiency benefits introduced by CogniGraph and the robustness of LiCoMemory for real-time interactive deployment.
Table 2: Detailed performance of LiCoMemory and baselines on LoCoMo in real-time interaction. $K_{G}$ stands for token consumption per session during context processing stage while $K_{R}$ stands for prompt token input. Accordingly, $T_{G}$ stands for latency per session of context processing stage $T_{R}$ stands for query latency.
Method Accuracy Token Latency
$K_{G}$ $K_{R}$ $T_{G}$ $T_{R}$
Zep 38.7% 212.5k 4.0k 2871s 5.71s
Mem0 54.68% 49.3k 2.2k 1772s 1.78s
Mem0g 55.82% 61.8k 2.4k 2081s 2.25s
A-Mem 44.12% 30.7k 4.1k 209s 4.63s
MemOS 54.08% 143.2k 2.6k 256s 2.71s
[…]
4.4 Ablation Study(RQ3)
We analyze the contribution of each major component of LiCoMemory through an ablation study, where individual modules are selectively removed to assess their impact across diverse evaluation scenarios. Figure 5 summarizes the results, showing that disabling different components leads to varying degrees of performance degradation, thereby revealing their complementary roles in supporting coherent long-term reasoning, temporal consistency, and effective cross-session retrieval.
Overall, all ablated variants exhibit clear performance drops, with the most severe degradation observed for w/o Structured retrieval. When retrieval is performed solely over extracted triples, ignoring session hierarchy and entity–relation structure, performance drops from 73.8/76.6 to 51.6/56.4 on the full dataset and from 64.7/73.5 to 40.6/48.2 in the multi-session setting, indicating fragmented retrieval and weakened factual grounding. Removing temporal weighting (w/o Temporal awareness) causes sharp declines in time-sensitive tasks, with temporal reasoning falling from 71.4/81.3 to 48.9/53.5 and knowledge update from 80.8/85.1 to 51.9/57.0, demonstrating the necessity of temporal signals for avoiding outdated evidence. Finally, disabling summary-level guidance (w/o Summary) consistently degrades performance (73.8/80.6 $\rightarrow$ 61.4/69.5 on the full dataset; 64.7/73.5 $\rightarrow$ 42.9/51.3 in multi-session settings), as retrieval becomes overly local and fails to capture higher-level contextual coherence. Together, these results confirm that structured retrieval, temporal awareness, and summary-level abstraction jointly underpin the robustness of LiCoMemory.